OpenStudy (anonymous):

Let f be twice differentiable with f (0) = 6, f (1) = 5, and f ′(1) = 2. Evaluate the integral...

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x f \prime \prime(x) dx\] which is \[f \prime(1) - f \prime(0)\] and my logic is that because there is an x in front of the prime, then \[f \prime(0) = 0\]

OpenStudy (anonymous):

am i in the right path?

OpenStudy (anonymous):

Not entirely - only partially (pun intended) You need integration by parts where u = x dv = f''

OpenStudy (anonymous):

i see...

OpenStudy (anonymous):

so i have \[x f \prime(x) - f \prime\]

OpenStudy (anonymous):

\[d(uv)=udv+vdu \rightarrow \int\limits d(uv)=uv=intudv+intvdu. \] Rearranging gives intudv=uv-intvdu

OpenStudy (anonymous):

\[ \int\limits udv=uv-\int\limits vdu\]

OpenStudy (anonymous):

oh poop, i made a mistake there

OpenStudy (anonymous):

y u dd

OpenStudy (anonymous):

so i would have \[x f \prime(x) - f(x) \]?

OpenStudy (anonymous):

xf' - VALUES_DIFFRENCE of f'

OpenStudy (anonymous):

VALUES_DIFFRENCE of xf' - VALUES_DIFFRENCE of f'

OpenStudy (anonymous):

is this correct? \[[1 f \prime(1) - f(1)] - [0 f \prime(0) - f(0)]\]

OpenStudy (anonymous):

or am i arranging them incorrectly?

OpenStudy (anonymous):

Second expr wrong change to f'(0} to f'(0)

OpenStudy (anonymous):

Change fourth to f'(0)

OpenStudy (anonymous):

how come? if \[dv= f \prime \prime\] then \[v= f \prime\] thus the antiderivative of \[v= f \prime\] would be f wouldn't it?

OpenStudy (amistre64):

u = x du = dx v = f' dv = f'' dx

OpenStudy (amistre64):

\[\int x~f''~dx=x~f'-\int f' dx\]

OpenStudy (anonymous):

so it is \[x f \prime(x) - f(x)\]

OpenStudy (amistre64):

yes, evaled at 0 and 1

OpenStudy (anonymous):

thank you, mikael was telling me that it was \[x f \prime - f \prime\] which didn't make sense to me

OpenStudy (amistre64):

\[(x f \prime(x) - f(x))^1-(x f \prime(x) - f(x))^0\] \[(1 f \prime(1) - f(1))-(0 f \prime(0) - f(0))\] \[f \prime(1) - f(1)+ f(0)\]

OpenStudy (anonymous):

awesome, thats what i was getting. thank you so much, you're the man... or woman ;)

OpenStudy (amistre64):

lol, good luck ;)